National Income Accounts.
X 100. Transfers 100. PIT 220. SST 60.
IBT1 40. NPI 40. I 100. CRE 20. C 700.
NFFI 100. Depreciation 60. IBT2 40. G 300. M 200.
Determine PI.
1
Expert's answer
2016-05-16T12:38:03-0400
The personal income is defined as follows:
PI= NI+transfers+NPI-CRE-SST-IBT2
Where
NI =GNP-depreciation-IBT1
GNP=GDP+NFFI=C+I+G+(X-M)+NFFI

GNP = 700 + 100 + 300 + (100-200) +100 =1100
NI =1100 – 60 – 40 =1000
PI = 1000 +100 + 40 – 20 – 60 – 40 = 1020